OpenStudy (ofmiceparade):

Find the horizontal or oblique asymptote of f(x) = 2 x squared plus 5 x plus 6, all over x plus 1.

OpenStudy (cheesecakekitten):

WILL MEDAL AND FAN!!! A gigabyte is a measure of a computer's storage capacity. One gigabyte holds about 1.0 x 109 bytes of information. If a company's computer network contains 3200 gigabytes of memory, how many bytes are in the network? A. 3.2 x 1010 B. 3.2 x 1012 C. 3.2 x 1011 D. 3.2 x 1013

OpenStudy (ofmiceparade):

@CheesecakeKitten Bruh..

OpenStudy (ofmiceparade):

@mathmale Hi

OpenStudy (mathmale):

OfMice: Hi. Could you please enter your problem statement symbolically (using either the Equation Editor or the Draw utility, below)? It's worth learning how to use those utilities. Find the horizontal or oblique asymptote of f(x) = 2 x squared plus 5 x plus 6, all over x plus 1. I'm unsure whether you mean (2x)^2 or 2x^2, for example.

OpenStudy (ofmiceparade):

@mathmale 2x^2+5x+6 ---------- x+1

OpenStudy (ofmiceparade):

@mathmale

OpenStudy (ofmiceparade):

@Kenshin hey

OpenStudy (kenshin):

@ofmiceparade Hi, do you understand what the question is asking of you?

OpenStudy (ofmiceparade):

@Kenshin Yea. I forgot how to do all this over winter break.

OpenStudy (phi):

you have an "improper" fraction. divide (x+1) into the top to make it a "mixed" fraction can you do that ?

OpenStudy (ofmiceparade):

? Im not exactly sure how to divide that to the top

OpenStudy (mathmale):

We are talking about "asymptotes" here, as specified by "horizontal or slant asymptote." OfMice: Before we proceed, please share what you think "asymptote" means in this context.

OpenStudy (phi):

it is hard to explain quickly if you don't know how (though I suspect you are supposed to know because that is how you solve this problem) see https://www.khanacademy.org/math/algebra2/arithmetic-with-polynomials/long-division-of-polynomials/v/polynomial-division for a quick review

OpenStudy (ofmiceparade):

An asymptote is a line or curve that approaches a given curve closely

OpenStudy (mathmale):

Actually, asymptotes are all straight lines, and graphs approach their asymptotes, rather than the other way around.

OpenStudy (ofmiceparade):

Oh

OpenStudy (mathmale):

What would be the equation of the horiz. asympt. of \[y=\frac{ x^2-4x + 5 }{ 2x^2 }?\]

OpenStudy (mathmale):

Have you found horiz. asy. before? If so, how?

OpenStudy (ofmiceparade):

...

OpenStudy (ofmiceparade):

I have no idea.

OpenStudy (ofmiceparade):

I haven't done this in over a month

OpenStudy (mathmale):

Have you thought of researching "long division" on the 'Net? or, if you have a textbook (which I hope you do), have you looked up "long div." there?

OpenStudy (mathmale):

I could explain this whole problem in a few minutes, but that would cheat you out of learning how to research info by yourself.

OpenStudy (ofmiceparade):

I just need help with question

OpenStudy (ofmiceparade):

I don't know where to start

OpenStudy (mathmale):

Suppose you let x grow larger and larger in the following: \[y=\frac{ x^2-4x + 5 }{ 2x^2 }\] As you do this, y will approach what value?

OpenStudy (ofmiceparade):

Division?

OpenStudy (ofmiceparade):

I'm clueless to what you just asked.

OpenStudy (kenshin):

When you have an improper fraction of algebraic expressions, after you manipulate it into a mixed fraction using long division, you will get a quotient and may get a remainder term afterwards which stays above the denominator. the quotient value is your horizontal or oblique asymptote, by horizontal it just means it is a 'y = constant' value, by oblique it means it's slanted like a diagonal line or curve in other cases (@mathmale which I think it can be a curve still not just line if say..you get y=x^2 as the quotient before the remainder)

OpenStudy (mathmale):

I needed to test your knowledge of "asymptotes." Without that understanding on your part, we'd have to start from scratch. You say you haven't done this type of work for a month or so...which implies that you had some familiarity with it in the past. What stops you from going back to review that material?

OpenStudy (mathmale):

Take another look at:\[y=\frac{ x^2-4x + 5 }{ 2x^2 }\] and drop some terms:\[y=\frac{ x^2}{ 2x^2 }\]

OpenStudy (ofmiceparade):

I'm trying to finish my first semester. This is one of the last assignments. I just want to finish.

OpenStudy (mathmale):

Now, what happens to y as x grows "very large?"

OpenStudy (ofmiceparade):

What do you even mean by that? I don't know. It changes?

OpenStudy (mathmale):

\[y=\frac{ x^2 }{ 2x^2 }\]

OpenStudy (mathmale):

sorry, my previous entry was very hard to read. This is what I meant (see above)./

OpenStudy (ofmiceparade):

Okay. Y = x^2/2x^2

OpenStudy (mathmale):

Yes. I'm asking you to determine what y value this expression approaches if x keeps growing larger and larger.

OpenStudy (ofmiceparade):

The y intercept?

OpenStudy (mathmale):

No, I'm asking you for a numerical limit. "As x grows larger without bound, y=(x^2) / (2x^2) approaches the value _______ "

OpenStudy (ofmiceparade):

y^2

OpenStudy (mathmale):

What is (x^2) / (x^2)?

OpenStudy (mathmale):

Asymptotes are usefun in graphing functions. For example, in the function we've been discussing, if x grows larger without bound, y approaches the limit (1/2). thus, the horizontal asymptote is a line described by y=1/2. Have you encountered this information / material before?

OpenStudy (ofmiceparade):

x times x over x times x

OpenStudy (mathmale):

^useful, not usefun.

OpenStudy (mathmale):

Yes, and what is x times x over x times x?

OpenStudy (ofmiceparade):

theres no number.. WHAT DO YOU MEAN

OpenStudy (ofmiceparade):

its an X

OpenStudy (mathmale):

this is algebra. x/x = 1. (x^2) / (x^2) = 1.

OpenStudy (mathmale):

What I am doing here is trying to introduce the concept of "asymptote." This concept seems to be either very new or just strange to y ou. If that's the case, this problem is not going to make any sense to you. I'm willing to spend time with you if you truly want to learn this material, but as a matter of principle, I won't be giving out answers. My goal is that you learn enough so that you can find your own answers properly.

OpenStudy (ofmiceparade):

oh

OpenStudy (ofmiceparade):

yea

OpenStudy (mathmale):

Here is the original question: "Find the horizontal or oblique asymptote of f(x) = 2 x squared plus 5 x plus 6, all over x plus 1." I'll go so far as to tell you that there is an asymptote here, and that it's a horiz. asympt., not an oblique (slant) asymptote. What form would the equation of a horiz. asymptote have? It's just a horiz. line.

OpenStudy (ofmiceparade):

y= something since theres no x point?

OpenStudy (mathmale):

\[\frac{ 2x^2+5x+6 }{ x+1 }\]

OpenStudy (ofmiceparade):

I already wrote that but now what?

OpenStudy (mathmale):

That "something" would be a numeric value. Look at the expression I've just typed in, above. the numerator has the order 2 (as in 2x^2) and the denominator has the order 1 (as in x+1). if you divide x+1 into 2x^2 + 5x + 6, you get the following: y becomes approximately equal to 2x + 3. This has the form y=mx + b and is therefore a straight line (asymptote line, slant asymptote).

OpenStudy (mathmale):

If you were to graph both the given function and this asymptote line, you would see that the curve approaches the line (like a horse appraoching a fence almost parallel to the fence).

OpenStudy (mathmale):

Do you have a graphing calculator? If so, would you be able to graph the function y=(2x^2+5x+6) / (x+1)?

OpenStudy (ofmiceparade):

Yea hang on, lemme get it

OpenStudy (ofmiceparade):

I graphed it

OpenStudy (mathmale):

Can you describe what your graph looks like? Is it a continuous curve or is it a curve with two separate parts?

OpenStudy (ofmiceparade):

Thats it

OpenStudy (mathmale):

Please see the following. I've used an online utility to graph the same function. My result agrees with yours. http://www.wolframalpha.com/input/?i=%282x%5E2%2B5x%2B6%29+%2F+%28x%2B1%29

OpenStudy (mathmale):

Now imagine a straight line drawn diagonally just above the left half of the graph and just below the right half of the graph. The curves practically touch the straight line as you move away from the origin. THAT is the "slant asymptote." Its equation is y=2x+3.

OpenStudy (ofmiceparade):

Oh Okay. Thanks so much!

OpenStudy (mathmale):

time to wrap this up. But first, any questions about what we've done here? I

OpenStudy (ofmiceparade):

@mathmale Nope. Thanks for spending like 30 minutes on me. I faned and medaled you

OpenStudy (mathmale):

I strongly encourage you to look up and review terms such as "asymptote,", "slant asympt.," horiz. asy. and vertical asy.

OpenStudy (mathmale):

...and you should know how to find them. thank you for your patience and perseverance and for the medal / fan. Take care!

OpenStudy (ofmiceparade):

Eh, im never using this in my real life. But thanks tho.

OpenStudy (mathmale):

Perhaps not, but learning HOW TO LEARN is extremely imkportant.

OpenStudy (ofmiceparade):

Its a bit more difficult when you don't care about the subject. I rather learn language

OpenStudy (mathmale):

I've forgotten at least 90% of the subject matter I learned in college. But I'm good at le3arning or re-learning stuff. That's what counts. Yes, I undrstand: you prefer verbal rather than math'l or scientific learning. That's OK. Just a preference. and greater talent in one area than in another. Best wishes to you.